<html>
  <head>

    <meta http-equiv="content-type" content="text/html; charset=utf-8">
  </head>
  <body bgcolor="#FFFFFF" text="#000000">
    Folks,<br>
    <br>
    Find the notes below and at<br>
    <br>
     
<a class="moz-txt-link-freetext" href="https://halldweb.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_July_27,_2015">https://halldweb.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_July_27,_2015</a><br>
    <br>
      -- Mark<br>
      ___________________<br>
    <br>
    <br>
    <div id="globalWrapper">
      <div id="column-content">
        <div id="content" class="mw-body" role="main"> <span dir="auto">Private:Hall
            D Staff Meeting, July 27, 2015</span><br>
          <div id="bodyContent" class="mw-body-content">
            <div id="siteSub"><br>
            </div>
            <div id="mw-content-text" dir="ltr" class="mw-content-ltr"
              lang="en">Present: Mary Jo Bailey, Fernando Barbosa,
              Eugene Chudakov (chair), Alexandre Deur, Sergey Furletov,
              Mark Ito, David Lawrence, Bert Manzlak, Paul Mattione,
              Lubomir Pentchev, Elton Smith, Sascha Somov, Justin
              Stevens, Simon Taylor, Glenn Young, Beni Zihlmann
              <br>
              <div id="toc" class="toc"><br>
                <br>
              </div>
              <span class="mw-headline" id="Eugene">Eugene</span><br>
              <ul>
                <li> PAC results came out</li>
                <li> solenoid review, two weeks ago
                  <ul>
                    <li> final report is out</li>
                    <li> thermostats should be looked at</li>
                    <li> lack of cryostability: cause not clear
                      <ul>
                        <li> local heating?</li>
                        <li> ice build up? in He volume? vacuum volume?</li>
                      </ul>
                    </li>
                    <li> He supply system will not see small
                      fluctuations
                      <ul>
                        <li> 20% variation might not be noticed</li>
                      </ul>
                    </li>
                    <li> pressure rise causes rise in resistance</li>
                  </ul>
                </li>
                <li> we need to make a space request for the DIRC to
                  Walt Akers
                  <ul>
                    <li> Justin: leave boxes at SLAC as long as possible</li>
                    <li> when they are moved here, move directly to Hall</li>
                    <li> will try to use one-bar-box read-out prototype
                      set-up at SLAC</li>
                    <li> will needs to devise tests to make sure boxes
                      made trip to JLab safely</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Mary_Jo">Mary Jo</span><br>
              <ul>
                <li> new proposal for hazard issues list
                  <ul>
                    <li> list reviewed annually</li>
                    <li> part of EH&S manual</li>
                    <li> old list revised: some items not strictly
                      issues, per se</li>
                    <li> rewritten from worker's point of view</li>
                    <li> linked to OSP system to ensure appropriate SME
                      review</li>
                    <li> methodology: ask questions for each area</li>
                    <li> comments to Mary Jo (<a class="moz-txt-link-abbreviated" href="mailto:mbailey@jlab.org">mbailey@jlab.org</a>)</li>
                    <li> proposed roll-out of new list: August 3, 2105</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Fernando">Fernando</span><br>
              <ul>
                <li> stuff is working</li>
                <li> current work: polarimeter
                  <ul>
                    <li> new enclosure, power distribution</li>
                  </ul>
                </li>
                <li> MO boards
                  <ul>
                    <li> ran without firmware in Spring, Ed working on
                      it</li>
                    <li> RF clock jitter: several fs, between crates:
                      150 ps</li>
                    <li> Ed J.: perhaps from loss of resolution at
                      rollover of F1-TDC</li>
                    <li> Ben: jitter tested out of clock distribution
                      system: 3 ps</li>
                    <li> in Hall there are TIs, STs</li>
                    <li> Paul and Elton looking at jitter</li>
                    <li> 4 crates involved: PS, TOF, tagger, FDC</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Eugene.2C_Engineering">Eugene,
                Engineering</span><br>
              <ul>
                <li> Tim working on acoustic detectors for solenoid</li>
                <li> work with SRF group</li>
                <li> Piezo-electric device attached to metal plate, from
                  there with rods to coils</li>
                <li> thermal events make no noise, motion of flux jumps
                  might</li>
              </ul>
              <br>
              <span class="mw-headline"
                id="Eugene.2C_End-of-Fiscal-Year_Purchases">Eugene,
                End-of-Fiscal-Year Purchases</span><br>
              <ul>
                <li> please send requests to Eugene within the week</li>
                <li> we have lots of TI boards</li>
                <li> we may be short of trigger distribution boards</li>
                <li> rapid access for Hall D requires CARM boxes, not
                  cheap</li>
              </ul>
              <br>
              <span class="mw-headline"
                id="Mark_and_David.2C_Git_Conversion">Mark and David,
                Git Conversion</span><br>
              <ul>
                <li> Git conversion generally going well
                  <ul>
                    <li> No major problems</li>
                    <li> We are climbing the learning curve</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="David">David</span><br>
              <ul>
                <li> sending a test set-up to UMass for CPP detector
                  tests
                  <ul>
                    <li> including an Flash-125</li>
                  </ul>
                </li>
                <li> working on Geant4, merging CPP with Richard Jones's
                  work on geometry</li>
              </ul>
              <br>
              <span class="mw-headline" id="Elton">Elton</span><br>
              <ul>
                <li> loan to UMass went quickly perhaps because CPP is
                  an approved experiment</li>
                <li> BCAL back on after power shutdown
                  <ul>
                    <li> HV channel tripped, on for one week now</li>
                    <li> humidity monitors were erratic, replaced
                      sensors, upstream and downstream about the same
                      now</li>
                  </ul>
                </li>
                <li> FCAL calibrations
                  <ul>
                    <li> looking at electrons 10% lower energy for
                      clusters, hits are included from out-of-time</li>
                    <li> see radial dependence in constants</li>
                  </ul>
                </li>
                <li> focus shifting to Monte Carlo, looking for limits
                  of hardware</li>
                <li> improvements to firmware for Flash-250s being
                  discussed
                  <ul>
                    <li> Sasha: document exists for scintillators</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Simon">Simon</span><br>
              <ul>
                <li> FCAL clusters including late hits, re-working
                  clusterizer adding timing cut</li>
                <li> CDC straws distorted, timing distributions
                  different on different sides of the wire, source of
                  long tails in timing dists
                  <ul>
                    <li> Mike Staib looking at cosmics, Simon at beam
                      data</li>
                    <li> bad straw fraction may be as high as 50%</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Sergey">Sergey</span><br>
              <ul>
                <li> looking at new FPGA code for F1TDCs</li>
                <li> asked for final version of F125 code from Cody
                  <ul>
                    <li> at least would like final version as far as DAQ
                      is concerned</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Paul">Paul</span><br>
              <ul>
                <li> submitted proposal to JSA for analysis workshop</li>
                <li> scaler gui's written, need test</li>
              </ul>
              <br>
              <span class="mw-headline" id="Mark">Mark</span><br>
              <ul>
                <li> Spring 2015 simulations: data generated</li>
                <li> Data Challenge 3: generating 10 TB of "raw" data</li>
              </ul>
              <br>
              <span class="mw-headline" id="Sascha">Sascha</span><br>
              <ul>
                <li> F1-TDC, Ed, read-out failure at high rate, Sergey
                  reproduced, Ed confirmed
                  <ul>
                    <li> new version of firmware coming</li>
                    <li> FIFO overflow causes FPGA code failure</li>
                  </ul>
                </li>
                <li> students are documenting, pair spectrometer
                  hodoscope, time calibration</li>
                <li> working on proposal for nuclear targets</li>
              </ul>
              <br>
              <span class="mw-headline" id="Alexandre">Alexandre</span><br>
              <ul>
                <li> looking at beamline data
                  <ul>
                    <li> no time coincidences between beam event and
                      magnet quench found</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Justin">Justin</span><br>
              <ul>
                <li> S&T starts tomorrow, runs for 3 days
                  <ul>
                    <li> has given practice talks, favorable reaction
                      from Management</li>
                  </ul>
                </li>
                <li> polarized rho's, fits being performed, spin density
                  matrix elements look do-able</li>
                <li> DIRC: prepare for review, writing TDR documents</li>
              </ul>
              <br>
              <span class="mw-headline" id="Lubomir">Lubomir</span><br>
              <ul>
                <li> FDC at design resolution with straight tracks
                  <ul>
                    <li> 80 micron wire position resolution with two
                      planes after correction for angle</li>
                    <li> 200 microns resolution from tracking residuals</li>
                  </ul>
                </li>
                <li> studies with Fe55 source
                  <ul>
                    <li> see effect of screening of wire</li>
                    <li> confirmed with Garfield simulation</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Beni">Beni</span><br>
              <ul>
                <li> FDC operational now, but not with normal gas
                  mixture
                  <ul>
                    <li> CO2 MFC died</li>
                    <li> CDC running off FDC system with 50-50 mix</li>
                    <li> 100% argon in CDC for a bit</li>
                  </ul>
                </li>
                <li> target retracted, upstream end of CDC accessible
                  <ul>
                    <li> Beni: work on grounding, noise on CDC even
                      after high frequency filters on HV lines, likely
                      noise from pre-amp cards</li>
                    <li> sees slope on baseline</li>
                    <li> need to talk to Tom if want to delay target
                      re-insertion</li>
                  </ul>
                </li>
                <li> working with FSU student Brad Cannon on TOF
                  calibration. Brad on-site for two weeks</li>
              </ul>
              <br>
              <span class="mw-headline" id="Bert">Bert</span><br>
              <ul>
                <li> 70% of all summer students at Lab working in
                  Physics
                  <ul>
                    <li> Rolf wants survey on student safety: labs,
                      halls, storage buildings</li>
                    <li> EH&S asking the questions</li>
                    <li> look at hazard awareness, work control
                      documents, required training, knowledge of limits
                      of work practices</li>
                    <li> supervisors responsibility nominally</li>
                  </ul>
                </li>
                <li> HDList being used, please continue to do so
                  <ul>
                    <li> see Bert if you need help</li>
                    <li> posting a printed copy in work area a good
                      practice</li>
                  </ul>
                </li>
                <li> about one week ago: golf cart event
                  <ul>
                    <li> accelerator got stuck, hit bumper stops</li>
                    <li> investigation on-going</li>
                    <li> test brakes if using golf carts</li>
                  </ul>
                </li>
              </ul>
              <br>
              <span class="mw-headline" id="Eugene.2C_Solenoid">Eugene,
                Solenoid</span><br>
              <ul>
                <li> planning test in two weeks</li>
                <li> test plan not firm</li>
                <li> an opportunity for test effect of field on
                  detectors, be prepared if this is a concern</li>
                <li> magnet on for 1 or 2 weeks</li>
                <li> considering re-doing cryostat on coil 4
                  <ul>
                    <li> only one not done</li>
                    <li> but coil 1 quenched first</li>
                  </ul>
                </li>
              </ul>
            </div>
            <div class="printfooter">
              Retrieved from "<a dir="ltr"
href="https://halldweb.jlab.org/wiki/index.php?title=Private:Hall_D_Staff_Meeting,_July_27,_2015&oldid=68916">https://halldweb.jlab.org/wiki/index.php?title=Private:Hall_D_Staff_Meeting,_July_27,_2015</a>"<br>
              <br>
            </div>
          </div>
        </div>
      </div>
    </div>
  </body>
</html>